OpenStudy (anonymous):

simplify; sec+csc / 1+tan

OpenStudy (mertsj):

\[\frac{\frac{1}{\cos }+\frac{1}{\sin }}{1+\frac{\sin }{\cos }}\times\frac{\sin \cos }{\sin \cos }=\frac{\sin +\cos }{\sin \cos +\sin ^2}=\frac{\sin +\cos }{\sin (\cos +\sin) }=\frac{1}{\sin }=\]

OpenStudy (mertsj):

csc

OpenStudy (anonymous):

thank you!!!!!

OpenStudy (mertsj):

yw